40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 27 21:48:31 2014, Jenne, Update, LSC, X green broadband PD not working???!? ALS_outofloop_27May2013_2.pdf
    Reply  Wed May 28 17:51:48 2014, manasa, Update, LSC, X green broadband PD NOT working 
       Reply  Wed May 28 19:15:38 2014, Koji, Update, LSC, X green broadband PD NOT working 
Message ID: 10000     Entry time: Wed May 28 17:51:48 2014     In reply to: 9996     Reply to this: 10001
Author: manasa 
Type: Update 
Category: LSC 
Subject: X green broadband PD NOT working 

Quote:

Grr.  I am very frustrated.  After lunch I redid alignment for both X and Y green systems (Yarm both at the end and on the PSL table, Xarm just on the PSL table).  After that realignment work, I cannot find a beatnote for the Xarm!!! 

At this point, I still hadn't touched anything on the X path (except the PZT input steering mirrors, remotely from the control room).  The beatnote was about the same size as it was on Friday, around -27dBm.  I went onto the PSL table and did the same alignment procedure that I had just done for the Yarm:  Remove the green trans PD and the accompanying lens so that I get far-field spots on the wall, and then steer the PSL green and the X green spots until they are nicely overlapped at both the camera (near-field) and on the wall.  I looked at the DC output of the beat PD, and centered the beam on the diode.  I put back the thorlabs DC transmission PD and the lens, and centered the beam on that.  However, after this work, I cannot find a beatnote for the X arm!  I still see the nice big Ygreen beatnote, and I have the PSL and Xend temperatures where they usually are (  abs(FSS Slow) < 0.1, and X end Slow around 10,090. )  I scanned -10,000 counts, and +5,000 counts from there, and still don't find a beatnote!

I went back inside, and I don't see an RF signal coming into the beatbox from the Xarm.  It's not the cable's fault though, since I then hooked the RF output of the beat PD to a 'scope, and still didn't see any beatnote.  The DC path of the PD is definitely seeing things, because when I switch the 'scope over to the DC output of the Xbeat PD, and I block/unblock the beam, I see the voltage step up and down as expected. 

I have not pulled out the Xgreen broadband PD, but unless someone else has a good idea of what to check, that might be one of the next things to do. 

Ideas of things I could try:

* Put the X broadband PD on the Y beatnote path to see if I see the same Y beatnote (use the port where the Y green trans PD is, since it has the coaligned beams, and a lens).

* Open the PD and see if anything on the RF path is fried.

* Move the Y PD over to the X path, to see if it sees the beatnote.

* ????

I made my attempts trying to figure out what was wrong.

Checking if we are at the right X end laser temperature: 
I aligned the arms and found the Y beatnote.I blocked the light falling on the X beat PD so that the RF analyser was only looking at the output from the Y beat PD. AT the RF analyser, I found the strong Y-PSL beatnote, the X-Y beat note and a weak  X-PSL beatnote. This confirmed that we have the X end laser at the right temperature to be able to detect the beatnote. Unblocking the light on the X beat PD did not bring in any additional peaks. 

Checking the RF cabling from the X beat PD to the beat box:
I swapped the RF cables such that the signal from the Y beat PD output was going to the X input on the beatbox. I could still see the beatnote on the RF analyser. This confirmed that there aren't any broken RF cables along the X path.

Checking X green PSL alignment:
I replaced the X beat PD with the Y beat PD to check if the alignment of X&PSL green are alright. I could find the X beat note this way without any alignment tweaking.

I suspect we probably have some RF component burnt in the X beat PD. Do we have any spares lying around? There is a Koji's box with a PD having the same serial number.

IFO status report for anyone who is looking to do some locking tonight : 
The Y beat PD is back along the Y path and I have confirmed the presence of Y-PSL beat note after replacing the PD.
The X beat PD has been removed and now rests on the electronics bench for checking. 

While aligning the arms today, I noticed that enabling LSC would cause misalignment of the ETMY suspension. I haven't tried to find out what has been causing this. Could be something similar to what was noticed with the ETMX suspension a couple of weeks ago elog9969 

ELOG V3.1.3-